OpenStudy (anonymous):

Write an equation of the line through each point. Use slope-intercept form (1,-1); parallel to y=2/5x-3

OpenStudy (anonymous):

y+1=2/5(x-1) y=2/5x-2/5-1 y=2/5x-7/5

OpenStudy (anonymous):

Can you break it down a little?

OpenStudy (anonymous):

slope intercept form is y-y1=m(x-x1) the slope from the first equation is 2/5 and because the lines are parallel they have the same slope so m=2/5 then i just plugged in the x and y values into the equation

OpenStudy (anonymous):

Thanks!